OpenStudy (anonymous):

What are the values of the functions y =x - 4 for x = 0, 1, 2 and 3? -4, -1, 2, 5 0 ,3 , 6, 9 -4, 2, 5, 8 0, 1, 5, 9

OpenStudy (anonymous):

y = 0 -4 y = 1 -4 y = 2 -4 y = 3 -4

OpenStudy (anonymous):

0-4 = -4 1-4 = -3 2-4 = -2 3-4 = -1

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

@larryboxaplenty

OpenStudy (anonymous):

@larryboxaplenty

OpenStudy (anonymous):

None of your answers has the set it should have. Did you type it wrong?

OpenStudy (anonymous):

Hol dn

OpenStudy (anonymous):

No it's right

OpenStudy (anonymous):

the set of answer is {-4, -3, -2, -1}, for y = x- 4, for x's 1,2,3,4